# On the metric of the jet bundle and similarity on weighted Dirichlet spaces

### v2

@@ Abstract.
In general, it is more difficult to formulate a sufficient condition for similarity than a necessary condition. We give sufficient conditions for a Cowen-Douglas operator with a positivity condition to be similar to the backward shift operator on weighted Dirichlet space. This condition involves the holomorphic jet bundle of the eigenvector bundle of the operator.
@@
Primary 47A45; Secondary 46E22, 46J20
The authors except for HK were supported by the National Natural Science Foundation of China, Grant No. 12371129 and 12471123. The second and the fifth authors appreciate the hospitality of the Department of Mathematics and Statistics at the University at Albany during their visit from Aug. 2023 to Feb. 2024.
@@ 1. Introduction
where $h_{\mathcal{E}_{T}}(w)=\left(\langle\gamma_{j}(w),\gamma_{i}(w)\rangle\right)_{i,j=1}^{n}$ denotes the metric associated with a holomorphic frame $\{\gamma_{1},\ldots,\gamma_{n}\}$ for $\mathcal{E}_{T}$. It is shown in [13] that the operators $T$ and $S$ in $\mathcal{B}_{n}(\Omega)$ are unitarily equivalent if and only if their complex bundles $\mathcal{E}_{T}$ and $\mathcal{E}_{S}$ are equivalent as Hermitian holomorphic vector bundles. They also proved that the curvature and its covariant derivatives form a complete set of unitary invariants. However, the similarity classification problem is much more complex and has not yet been fully resolved.
@@
For operators $T$ and $S$ in $\mathcal{B}_{1}(\mathbb{D}),$ each having $\overline{\mathbb{D}}$ as a $K$-spectral set, the Cowen-Douglas conjecture ([13], §$\S$4.35) states that $T$ is similar to $S$ if and only if $\lim\limits_{|w|\rightarrow 1}\mathcal{K}_{T}(w)/\mathcal{K}_{S}(w)=1,$ where $\mathcal{K}_{T}$ denotes the curvature of $\mathcal{E}_{T}$. Clark and Misra gave a positive result in [10], confirming that the conjecture is sufficient for the class of Toeplitz operators considered in [12]. Subsequently, they in [11] introduced a weighted backward shift operator to show that the conjecture is unnecessary, and found that the ratio of the metrics associated with Hermitian holomorphic line bundles is a better similarity invariant than the ratio of the curvatures, which is the main point of the note. In addition, Clark and Misra showed that the backward shift operator $T$, with $\|T\|\leq 1$, is similar to the backward shift operator $S_{\alpha}$ with weight sequence $\left\{[(n+1)/(n+2)]^{\alpha/2}\right\}_{n\geq 0}$ if and only if the ratio $a_{w}:=\frac{h_{\mathcal{E}_{T}}(w)}{h_{\mathcal{E}_{S_{\alpha}}}(w)}$ of the metric is bounded and bounded from 00 ([11], Theorem 1). In the third section of this note, we will use the ratio of metrics and the operator realization of holomorphic vector bundles to provide necessary and sufficient conditions for the similarity of the backward shift operators on the Hardy space and the weighted Bergman space.
@@
There have been some new developments in the past four decades on the similarity problem for the weighted backward shift operators on the spaces $\mathcal{H}_{k}$ [15, 26, 27, 30, 33, 34, 39, 42, 57]. Uchiyama used a holomorphic frame of Hermitian holomorphic vector bundles to show the similarity between a contraction and the backward shift operator on the Hardy space [55]. Douglas, Treil and the third author gave a necessary and sufficient condition for an $n$-hypercontraction to be similar to the backward shift operator on a weighted Bergman spaces by using the trace of the curvature [16, 38]. Sarkar and the first author described the similarity of backward shift operators on weighted Bergman spaces using Hilbert modules [29]. The third author gave a necessary condition for the similarity of the backward shift on the Dirichlet space [37]. Note that the original Cowen-Douglas conjecture does not include the backward shift operator in the Dirichlet space and operators similar to it, see Remark 2 in [11]. In this note, we introduce holomorphic jet bundle, which is closely related to holomorphic vector bundles, to consider the similarity of the backward shift operators on the weighted Dirichlet spaces.
@@
In [40], Lin considered the operator realization problem of when a function $f\in H^{\infty}(\mathbb{D})$ gives rise to a Cowen-Douglas operator and proved that the tensor product $\mathcal{E}_{T}\otimes\mathcal{E}_{S}$ from operators $T\in\mathcal{B}_{m}(\mathbb{D})$ and $S\in\mathcal{B}_{n}(\mathbb{D})$ is indeed a Hermitian holomorphic eigenvector bundle over $\mathbb{D}$ of some Cowen-Douglas operator in $\mathcal{B}_{nm}(\mathbb{D})$. We are interested in using the geometric quantities of the Hermitian holomorphic eigenvector bundle associated with a Cowen-Douglas operator to characterize similarity. In particular, we will use the metric of the Hermitian holomorphic eigenvector bundle and associated holomorphic jet bundles to give a similarity criteria.
The paper is organized as following. Section 2 fixes the notation and executes some preliminaries. Section 3 deals with some extension results, Theorem 3.4 and Proposition 3.6, which are given in terms of the tensor product of holomorphic vector bundles and the holomorphic vector bundles induced by the defect operator corresponding to the $n$-hypercontraction. Section 4 shows two sufficient conditions for operator $T\in\mathcal{L}(\mathcal{H})$ to be similar to the backward shift operators $D^{*}_{\alpha}$ on the weighted Dirichlet space $\mathcal{D}_{\alpha}$, namely Theorem 4.3 and Theorem 4.9, the techniques used for these results are the jet bundles of holomorphic complex bundles and the properties of multiplier algebras in weighted Dirichlet spaces.
@@ 2.1. Multipliers
It is clear now that any multiplier $\Phi$ is bounded and holomorphic, that is, $\Phi\in H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$, where $H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$ is the space of bounded analytic functions defined on $\mathbb{D}$ whose function values are bounded linear operators from a Hilbert space $E_{1}$ to another one $E_{2}$. When $\text{Mult}(\mathcal{H}_{K})=H^{\infty}(\mathbb{D})$, then $\text{Mult}(\mathcal{H}\otimes E_{1},\mathcal{H}_{K}\otimes E_{2})=H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$.
Since a bounded linear operator on the Hardy space commuting with the operator of multiplication by $z$ is given by the multiplication operator by a function $\varphi\in H^{\infty}(\mathbb{D})$ with norm $\|\varphi\|_{\infty},$ the corona problem is closely related to the multipliers on the reproducing kernel Hilbert space of analytic functions (see [7, 56]). The original corona problem posed by Kakutani in 1941 asks whether the unit disk $\mathbb{D}$ is dense in the maximal ideal space of $H^{\infty}(\mathbb{D})$. The question was later answered in the affirmative by Carleson in [9]. In [23], Hörmander introduced a new and a much simpler method for solving the problem which was popularized by Wolff (see [20, 45]). There have been many versions of the operator corona theorem given by Fuhrmann, Rosenblum, Tolokonnikov, Uchiyama, and Vasyunin [19, 46, 53, 54]. In particular, we have the following result by Fuhrmann [53]:
@@ 2.2. Model theorem
The original model theorem states that a contraction $T\in\mathcal{L}(\mathcal{H})$ with $\lim\limits_{j\rightarrow\infty}\|T^{j}h\|=0$ for all $h\in\mathcal{H}$ is unitarily equivalent to the backward shift operator restricted to some co-invariant subspace of a vector-valued Hardy space [18, 45, 48, 51]. Numerous analogues have been proposed in [1, 2, 4, 5, 6, 43, 44]. In particular, we mention the model involving the backward shift on the Hardy space, weighted Bergman space, and weighted Dirichlet space given by Sz.-Nagy-Foiaş [51], Agler [1, 2], and Müller [43], respectively. The result has been previously used by Douglas, Treil and the third author to give a necessary and sufficient condition for the similarity to the backward shift on the Hardy space [38] and the weighted Bergman space [16]. Subsequently, the third author to give a necessary condition for the similarity to the backward shift on the Dirichlet space [37].
@@
We denote $\mathcal{M}_{n}$ as a Hilbert space with the reproducing kernel $K_{n}(z,w)=\frac{1}{(1-\overline{w}z)^{n}}$ for $z,w\in\mathbb{D}$, where $n$ is a positive integer. As can be easily checked, different function spaces correspond to different $n$’s: the Hardy space for $n=1$ and the weighted Bergman spaces for $n\geq 2$. The vector-valued space $\mathcal{M}_{n,E}$ for a separable Hilbert space $E$ can be defined as previously. The forward shift $S_{n,E}$ on $\mathcal{M}_{n,E}$ is defined as $S_{n,E}f(z)=zf(z),$ and the backward shift operator $S^{*}_{n,E}$ is its adjoint. We are ready to state the following theorem by Agler [1, 2].
@@
For general unilateral shift operators, the work of Shields given in [49] is quite extensive. The result that we need the most in this note is the following, which easily shows that the backward shift operators on the Hardy space, the weighted Bergman spaces, and the weighted Dirichlet spaces are not similar.
@@
Let $T_{1}$ and $T_{2}$ be unilateral weighted shifts with weight sequences $\{\alpha_{i}\}_{i\geq 1}$ and $\{\beta_{i}\}_{i\geq 1}$, respectively. Then $T_{1}\sim_{s}T_{2}$ if and only if there exist constants $C_{1}$ and $C_{2}$ such that for every positive integer $l$,
$0<C_{1}\leq\left|\frac{\alpha_{k}\alpha_{k+1}\cdots\alpha_{l}}{\beta_{k}\beta_{k+1}\cdots\beta_{l}}\right|\leq C_{2}$
@@ 3. Similarity in the class $\mathcal{B}_{m}(\mathbb{D})$
$(2)\Rightarrow(1)$: Since $0<C_{1}I_{n}\leq h(w)\leq C_{2}I_{n},$
$C_{1}\leq\lambda_{i}(w)\leq C_{2}\quad\text{ and}\quad 0<nC_{1}\leq\sum\limits_{i=1}^{n}\|\gamma_{i}(w)\|^{2}=\sum\limits_{i=1}^{n}\lambda_{i}(w)\leq nC_{2},$
so that the function $F\in H^{\infty}_{\mathbb{C}^{n}\rightarrow E}(\mathbb{D})$ satisfies
$F(w)^{*}F(w)=\left(\langle\gamma_{j}(w),\gamma_{i}(w)\rangle\right)_{i,j=1}^{n}\geq C_{1},\quad w\in\mathbb{D}.$
@@
For the Cowen-Douglas conjecture, Clark and Misra noted in [11] that the ratio $a_{w}$ of the metrics is more suitable as a similarity invariant for the Cowen-Douglas operator. For operators $T\in\mathcal{B}_{m}(\mathbb{D})$ and $S\in\mathcal{B}_{m}(\mathbb{D})$, if $\mathcal{E}_{T}=\mathcal{E}_{S}\otimes\mathcal{E}$ for some Hermitian holomorphic vector bundle $\mathcal{E}$ over $\mathbb{D}$, then the metrics of these holomorphic vector bundles satisfy $h_{\mathcal{E}_{T}}=h_{\mathcal{E}_{S}}\otimes h_{\mathcal{E}}.$ In particular, when $\mathcal{E}_{T}$ and $\mathcal{E}_{S}$ are holomorphic line bundles, the ratio of their metrics is the metric of holomorphic line bundle $\mathcal{E}$, and thus $a_{w}:=h_{\mathcal{E}}(w)=\frac{h_{\mathcal{E}_{T}}(w)}{h_{\mathcal{E}_{S}}(w)}$. From the model theorems given in [1, 2, 4, 43, 44, 45, 52] and the operator theoretic realization in [40], we know many operators whose associated holomorphic eigenvector bundles possess a tensor structure. Now provide detailed proof of the Cowen-Douglas operator with $n$-hypercontraction possessing this property, which was proposed in [2].
@@
$Vx=\sum\limits_{k=0}^{\infty}\frac{z^{k}}{\|z^{k}\|_{\mathcal{M}_{n}}^{2}}\otimes D_{T}T^{k}x$
@@
and the eigenvector bundle $\ker(S^{*}_{n,E}|_{\mathcal{N}}-w)=\bigvee\{K_{n}(\cdot,\overline{w})D_{T}\gamma_{i}(w):1\leq i\leq m\}.$ From the eigenvector subspaces
@@
are analytically dependent on $w\in\mathbb{D}$, we can say the same for the subspace $\mathcal{E}_{D_{T}}(w)=\bigvee\{D_{T}\gamma_{i}(w):1\leq i\leq m\}$. So $\mathcal{E}_{D_{T}}=\{(w,x)\in\mathbb{D}\times\mathcal{E}(w):x\in\mathcal{E}_{D_{T}}(w)\}$ is a $m$-dimensional Hermitian holomorphic vector bundle over $\mathbb{D}$ with the natural projection $\pi,\pi(w,x)=w,$ and $\{D_{T}\gamma_{1},\ldots,D_{T}\gamma_{m}\}$ is a holomorphic frame of $\mathcal{E}_{D_{T}}$. Hence, $\mathcal{E}_{T}$ and $\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$ are equivalent as Hermitian holomorphic vector bundles. ∎
For the $n$-hypercontraction $T\in\mathcal{B}_{m}(\mathbb{D})$, let $\mathcal{E}_{D_{T}}$ be the holomorphic complex bundle induced by the defect operator $D_{T}$ of $T$, which has the following definition.
@@
If $T\in\mathcal{B}_{m}(\mathbb{D})$ is an $n$-hypercontraction and $\gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ is a holomorphic frame of $\mathcal{E}_{T}$, we said $\mathcal{E}_{D_{T}}=\{(w,x)\in\mathbb{D}\times\mathcal{E}_{D_{T}}(w):x\in\mathcal{E}_{D_{T}}(w)\}$ is a Hermitian holomorphic vector bundle over $\mathbb{D}$ of rank $m$ induced by the defect operator $D_{T}$ of $T$, where $\mathcal{E}_{D_{T}}(w)=\bigvee\{D_{T}\gamma_{i}(w):1\leq i\leq m\}$.
For the $n$-hypercontraction in $\mathcal{B}_{1}(\mathbb{D})$, Theorem 1 in [11] and Theorem 3.4 in [25] show that the ratio of the metrics of holomorphic line bundles is a similarity invariant for the backward shift operators on the Hardy space and the weighted Bergman spaces. More generally, for the similarity of operators in $\mathcal{B}_{m}(\mathbb{D})$, we use the multiplier algebra and Lemma 3.2 to give the following theorem, the main theorem of this section, concerns a similarity criteria for $n$-hypercontraction $T\in\mathcal{B}_{m}(\mathbb{D})$ by using the metric $h_{\mathcal{E}_{D_{T}}}$ of the holomorphic complex bundle $\mathcal{E}_{D_{T}}$:
@@
Let $T\in\mathcal{B}_{m}(\mathbb{D})$ be an $n$-hypercontraction, then the following statements are equivalent:
@@
    There exist constants $C_{1}$ and $C_{2}$ such that the metric matrix $h_{D_{T}}(w)$ of $\mathcal{E}_{D_{T}}$ in some holomorphic frame satisfies
@@
    There exists a holomorphic frame $\widetilde{\gamma}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{m}\}$ of $\mathcal{E}_{D_{T}}$ such that each $\|\widetilde{\gamma}_{i}(w)\|$ is uniformly bounded on $\mathbb{D}$, and there are constants $A_{1}$ and $A_{2}$ such that the metric matrix $h_{D_{T}}(w)$ under holomorphic frame $\widetilde{\gamma}$ satisfies
    $0<A_{1}\leq\det h_{\mathcal{E}_{D_{T}}}(w)\leq A_{2}\quad w\in\mathbb{D}.$
@@
$G(w)F(w)=I_{\mathbb{C}^{m}}\quad w\in\mathbb{D}.$
@@
Considering the composite of the unitary operator $V$, shown Lemma 3.2, with $X$, we see that the invertible operator $X^{-1}V$ satisfies $(X^{-1}V)\mathcal{E}_{T}(w)=\mathcal{E}_{S^{*}_{n,\mathbb{C}^{m}}}(w)$ for all $w\in\mathbb{D}$. It follows that $T$ is similar to $S^{*}_{n,\mathbb{C}^{m}}$.

### v3

# On the metric of the jet bundle and similarity of Cowen-Douglas operators

@@ Abstract.
The study of Cowen-Douglas operators not only involves traditional operator-theoretic tools but also concepts and results from complex geometry on holomorphic vector bundles. We make use of the ratio of the metric matrices first considered by Clark and Misra and a model theorem by Agler to describe the similarity of backward shift operators on analytic function spaces whose multiplier algebras are the space of bounded analytic functions. It is well-known that, in general, it becomes much more complicated to formulate a sufficient condition for similarity than a necessary one. We also give a sufficient condition for a Cowen-Douglas operator to be similar to the backward shift operator on the Dirichlet space with weights by introducing a condition on the jet bundle of a holomorphic vector bundle. Note that the multiplier algebras of these spaces do not coincide with the space of bounded, analytic functions as in other analytic functions spaces, requiring a different approach. The results by Müller on operator models related to Dirichlet shifts and by Kidane and Trent on the corona problem for the multiplier algebras of weighted Dirichlet spaces are indispensable tools in attaining this similarity result.
@@
Primary 47B13; Secondary 47A45, 32L05, 51M15, 53C07
The authors except for the third author were supported by the National Natural Science Foundation of China, Grant No. 12371129 and 12471123, and the Hebei Natural Science Foundation, Grant No. A2023205045. The second and the fourth authors appreciate the hospitality of the Department of Mathematics and Statistics at the University at Albany during their visit from Aug. 2023 to Feb. 2024.
@@ 1. Introduction
where $h_{\mathcal{E}_{T}}(w)=\left(\langle\gamma_{j}(w),\gamma_{i}(w)\rangle\right)_{i,j=1}^{n}$ denotes the metric associated with a holomorphic frame $\{\gamma_{1},\ldots,\gamma_{n}\}$ for $\mathcal{E}_{T}$. It is shown in [13] that operators $T$ and $S$ in $\mathcal{B}_{n}(\Omega)$ are unitarily equivalent if and only if their complex bundles $\mathcal{E}_{T}$ and $\mathcal{E}_{S}$ are equivalent ($\mathcal{E}_{S}\sim_{u}\mathcal{E}_{T}$) as Hermitian holomorphic vector bundles, that is, there exists a unitary operator $U\in\mathcal{L}(\mathcal{H})$ such that $\mathcal{E}_{S}(w)=U(\mathcal{E}_{T}(w))$ for all $w\in\Omega$. They also proved that the curvature and its covariant derivatives form a complete set of unitary invariants. However, the similarity classification problem is much more complex and has not yet been fully resolved.
@@
For operators $T$ and $S$ in $\mathcal{B}_{1}(\mathbb{D}),$ each having $\overline{\mathbb{D}}$ as a $K$-spectral set, the Cowen-Douglas conjecture ([13], §$\S$4.35) states that $T$ is similar to $S$ if and only if $\lim\limits_{|w|\rightarrow 1}\mathcal{K}_{\mathcal{E}_{T}}(w)/\mathcal{K}_{\mathcal{E}_{S}}(w)=1$. Clark and Misra gave a positive result in [10], confirming that the conjecture is sufficient for the class of Toeplitz operators considered in [12]. Subsequently, they in [11] introduced a weighted backward shift operator to show that the conjecture is not necessary, and found that the ratio of the metrics associated with Hermitian holomorphic line bundles is a better similarity invariant than the ratio of the curvatures, the main point of the note. In addition, Clark and Misra showed that the backward shift operator $T$, with $\|T\|\leq 1$, is similar to the backward shift operator $S_{\alpha}$ with weight sequence $\left\{[(n+1)/(n+2)]^{\alpha/2}\right\}_{n\geq 0}$ if and only if the ratio $a_{w}:=\frac{h_{\mathcal{E}_{T}}(w)}{h_{\mathcal{E}_{S_{\alpha}}}(w)}$ of the metric is bounded and bounded away from 00 ([11], Theorem 1). In the third section of this paper, we will use the ratio of the metrics and the operator realization of holomorphic vector bundles to provide necessary and sufficient conditions for the similarity of the backward shift operators on the Hardy and weighted Bergman spaces.
@@
There have been some new developments in the past four decades on the similarity problem for the weighted backward shift operators on the spaces $\mathcal{H}_{k}$ [15, 26, 27, 30, 33, 34, 39, 42, 57]. Uchiyama used a holomorphic frame of Hermitian holomorphic vector bundles to show the similarity between a contraction and the backward shift operator on the Hardy space [55]. Douglas, Treil and the third author gave a necessary and sufficient condition for an $n$-hypercontraction to be similar to the backward shift operator on weighted Bergman spaces by using the trace of the curvature [16, 38]. Sarkar and the first author described the similarity of backward shift operators on weighted Bergman spaces using Hilbert modules [29]. The third author gave a necessary condition for the similarity of the backward shift on the Dirichlet space [37]. Note that the original Cowen-Douglas conjecture does not include the backward shift operator in the Dirichlet space and operators similar to it, see Remark 2 in [11]. In this paper, we introduce the holomorphic jet bundle in the description, which is closely related to holomorphic vector bundles, to consider the similarity of the backward shift operators on weighted Dirichlet spaces.
@@
In [40], Lin considered the operator realization problem of when a function $f\in H^{\infty}(\mathbb{D})$ gives rise to a Cowen-Douglas operator and proved that the tensor product $\mathcal{E}_{T}\otimes\mathcal{E}_{S}$ from operators $T\in\mathcal{B}_{m}(\mathbb{D})$ and $S\in\mathcal{B}_{n}(\mathbb{D})$ is indeed a Hermitian holomorphic eigenvector bundle over $\mathbb{D}$ of some Cowen-Douglas operator in $\mathcal{B}_{nm}(\mathbb{D})$. We are interested in using geometric entities of the Hermitian holomorphic eigenvector bundle associated with a Cowen-Douglas operator to characterize similarity. In particular, we will use the metric of the Hermitian holomorphic eigenvector bundle and the associated holomorphic jet bundles to give a similarity criteria.
The paper is organized as follows: in Section 2, we introduce some necessary notations and present preliminary material. Section 3 deals with some extension results, Theorem 3.4 and Proposition 3.6, which are given in terms of the tensor product of holomorphic vector bundles and the holomorphic vector bundles induced by the defect operator corresponding to an $n$-hypercontraction. Section 4 includes two sufficient conditions for the operator $T\in\mathcal{L}(\mathcal{H})$ to be similar to the backward shift operators $D^{*}_{\alpha}$ on weighted Dirichlet spaces $\mathcal{D}_{\alpha}$, namely Theorems 4.3 and 4.9; the techniques used for these results involve the jet bundles of holomorphic complex bundles and properties of multiplier algebras of weighted Dirichlet spaces.
@@ 2.1. Multipliers
It is clear now that any multiplier $\Phi$ is bounded and holomorphic, that is, $\Phi\in H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$, where $H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$ is the space of bounded analytic functions defined on $\mathbb{D}$ whose function values are bounded linear operators from a Hilbert space $E_{1}$ to another one $E_{2}$. When $\text{Mult}(\mathcal{H}_{K})=H^{\infty}(\mathbb{D})$, then $\text{Mult}(\mathcal{H}_{K}\otimes E_{1},\mathcal{H}_{K}\otimes E_{2})=H^{\infty}_{E_{1}\rightarrow E_{2}}(\mathbb{D})$.
Since a bounded linear operator on the Hardy space $H^{2}(\mathbb{D})$ commuting with the operator of multiplication by $z$ is given by the multiplication operator by a function $\varphi\in H^{\infty}(\mathbb{D})$ with norm $\|\varphi\|_{\infty},$ the corona problem is closely related to the multipliers on reproducing kernel Hilbert spaces of analytic functions (see [7, 56]). The original corona problem posed by Kakutani in 1941 asks whether the unit disk $\mathbb{D}$ is dense in the maximal ideal space of $H^{\infty}(\mathbb{D})$. The question was later answered in the affirmative by Carleson in [9]. In [23], Hörmander introduced a new and a much simpler method for solving the problem which was popularized by Wolff (see [20, 45]). There have been many versions of the operator corona theorem given by Fuhrmann, Rosenblum, Tolokonnikov, Uchiyama, and Vasyunin [19, 46, 53, 54]. In particular, we have the following result by Vasyunin [53]:
@@ 2.2. Model theorem
The original model theorem states that a contraction $T\in\mathcal{L}(\mathcal{H})$ with $\lim\limits_{j\rightarrow\infty}\|T^{j}h\|=0$ for all $h\in\mathcal{H}$ is unitarily equivalent to the backward shift operator restricted to some co-invariant subspace of a vector-valued Hardy space [18, 45, 48, 51]. Numerous analogues have been proposed in [1, 2, 4, 5, 6, 43, 44]. In particular, we mention the models involving the backward shift on the Hardy space, weighted Bergman space, and weighted Dirichlet space given by Sz.-Nagy-Foiaş [51], Agler [1, 2], and Müller [43], respectively. The results have been previously used by Douglas, Treil and the third author to describe similarity to the backward shift on these spaces in [16], [37], and [38].
@@
For a positive integer $n$, we let $\mathcal{M}_{n}$ be the Hilbert space with reproducing kernel $K_{n}(z,w)=\frac{1}{(1-\overline{w}z)^{n}}$ for $z,w\in\mathbb{D}$. As can be easily checked, different function spaces correspond to different $n$’s: the Hardy space for $n=1$ and the weighted Bergman spaces for $n\geq 2$. The vector-valued space $\mathcal{M}_{n,E}$ for a separable Hilbert space $E$ can also be defined. The forward shift $S_{n,E}$ on $\mathcal{M}_{n,E}$ is defined as $S_{n,E}f(z)=zf(z),$ and the backward shift operator $S^{*}_{n,E}$ is its adjoint. We are ready to state the following theorem by Agler [1, 2].
@@
For general unilateral shift operators, the work of Shields given in [49] is quite extensive. The result that we need the most in this paper is the next theorem, which can also be used to give an easy proof that the backward shift operators on the Hardy space, the weighted Bergman spaces, and the weighted Dirichlet spaces are not similar to each other.
@@
Let $T_{1}$ and $T_{2}$ be unilateral weighted shifts with non-zero weight sequences $\{\alpha_{i}\}_{i\geq 1}$ and $\{\beta_{i}\}_{i\geq 1}$, respectively. Then $T_{1}\sim_{s}T_{2}$ if and only if there exist constants $C_{1}$ and $C_{2}$ such that for every positive integer $l$,
$0<C_{1}\leq\left|\frac{\alpha_{k}\alpha_{k+1}\cdots\alpha_{l}}{\beta_{k}\beta_{k+1}\cdots\beta_{l}}\right|\leq C_{2},$
@@ 3. Similarity in the class $\mathcal{B}_{m}(\mathbb{D})$
$(2)\Rightarrow(1)$: Since $0<C_{1}I_{n}\leq h(w)\leq C_{2}I_{n},$ the function $F\in H^{\infty}_{\mathbb{C}^{n}\rightarrow E}(\mathbb{D})$ satisfies
$F(w)^{*}F(w)=\left(\langle\gamma_{j}(w),\gamma_{i}(w)\rangle\right)_{i,j=1}^{n}=h(w)\geq C_{1}I_{n},\quad w\in\mathbb{D}.$
@@
Regarding the Cowen-Douglas conjecture, Clark and Misra noted in [11] that the ratio $a_{w}$ of the metrics is more suitable as a similarity invariant for Cowen-Douglas operators. For operators $T\in\mathcal{B}_{m}(\mathbb{D})$ and $S\in\mathcal{B}_{m}(\mathbb{D})$, if $\mathcal{E}_{T}=\mathcal{E}_{S}\otimes\mathcal{E}$ for some Hermitian holomorphic vector bundle $\mathcal{E}$ over $\mathbb{D}$, then the metrics of these holomorphic vector bundles satisfy $h_{\mathcal{E}_{T}}=h_{\mathcal{E}_{S}}\otimes h_{\mathcal{E}}.$ In particular, when $\mathcal{E}_{T}$ and $\mathcal{E}_{S}$ are holomorphic line bundles, the ratio of their metrics is the metric of holomorphic line bundle $\mathcal{E}$, and thus $a_{w}:=h_{\mathcal{E}}(w)=\frac{h_{\mathcal{E}_{T}}(w)}{h_{\mathcal{E}_{S}}(w)}$. From the model theorems given in [1, 2, 4, 43, 44, 45, 52] and the operator theoretic realization in [40], we know of many operators whose associated holomorphic eigenvector bundles possess a tensor structure. We now provide details on how to construct an $n$-hypercontractive Cowen-Douglas operator with this property, using the result of Agler in [2].
@@
$Vx=\sum\limits_{k=0}^{\infty}\frac{z^{k}}{\|z^{k}\|_{\mathcal{M}_{n}}^{2}}\otimes D_{T}T^{k}x,$
@@
and the eigenvector bundle $\ker(S^{*}_{n,E}|_{\mathcal{N}}-w)$ can be expressed as
$\ker(S^{*}_{n,E}|_{\mathcal{N}}-w)=\bigvee\{K_{n}(\cdot,\overline{w})\otimes D_{T}\gamma_{i}(w):1\leq i\leq m\}.$
Since the eigenspaces
@@
depend analytically on $w\in\mathbb{D}$, we can say the same for the subspace $\mathcal{E}_{D_{T}}(w)=\bigvee\{D_{T}\gamma_{i}(w):1\leq i\leq m\}$. Therefore, $\mathcal{E}_{D_{T}}=\{(w,x)\in\mathbb{D}\times\mathcal{E}(w):x\in\mathcal{E}_{D_{T}}(w)\}$ is an $m$-dimensional Hermitian holomorphic vector bundle over $\mathbb{D}$ with the natural projection $\pi,\pi(w,x)=w,$ and $\{D_{T}\gamma_{1},\ldots,D_{T}\gamma_{m}\}$ is a holomorphic frame of $\mathcal{E}_{D_{T}}$. Hence, $\mathcal{E}_{T}$ and $\mathcal{E}_{S^{*}_{n}}\otimes\mathcal{E}_{D_{T}}$ are equivalent as Hermitian holomorphic vector bundles. ∎
For an $n$-hypercontraction $T\in\mathcal{B}_{m}(\mathbb{D})$, let $\mathcal{E}_{D_{T}}$ be the holomorphic complex bundle induced by the defect operator $D_{T}$ of $T$. The precise definition is as follows:
@@
If $T\in\mathcal{B}_{m}(\mathbb{D})$ is an $n$-hypercontraction and $\gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ is a holomorphic frame of $\mathcal{E}_{T}$, we say that $\mathcal{E}_{D_{T}}=\{(w,x)\in\mathbb{D}\times\mathcal{E}_{D_{T}}(w):x\in\mathcal{E}_{D_{T}}(w)\}$ is a Hermitian holomorphic vector bundle over $\mathbb{D}$ of rank $m$ induced by the defect operator $D_{T}$ of $T$, where $\mathcal{E}_{D_{T}}(w)=\bigvee\{D_{T}\gamma_{i}(w):1\leq i\leq m\}$.
For an $n$-hypercontraction in $\mathcal{B}_{1}(\mathbb{D})$, Theorem 1 in [11] and Theorem 3.4 in [25] show that the ratio of the metrics of holomorphic line bundles is a similarity invariant for the backward shift operators on the Hardy and weighted Bergman spaces. More generally, for the similarity of operators in $\mathcal{B}_{m}(\mathbb{D})$, we use the multiplier algebra and Lemma 3.2 to give the following theorem, the main theorem of this section, that provides a similarity criteria for an $n$-hypercontraction $T\in\mathcal{B}_{m}(\mathbb{D})$ in terms of the metric $h_{\mathcal{E}_{D_{T}}}$ of the holomorphic complex bundle $\mathcal{E}_{D_{T}}$:
@@
For an $n$-hypercontraction $T\in\mathcal{B}_{m}(\mathbb{D})$, the following statements are equivalent:
@@
    There exist constants $C_{1}$ and $C_{2}$ such that the metric matrix $h_{\mathcal{E}_{D_{T}}}(w)$ of $\mathcal{E}_{D_{T}}$ in some holomorphic frame satisfies
@@
    There exists a holomorphic frame $\widetilde{\gamma}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{m}\}$ of $\mathcal{E}_{D_{T}}$ such that each $\|\widetilde{\gamma}_{i}(w)\|$ is uniformly bounded on $\mathbb{D}$, and that the metric matrix $h_{\mathcal{E}_{D_{T}}}(w)$ under $\widetilde{\gamma}$ satisfies
    $0<A_{1}\leq\det h_{\mathcal{E}_{D_{T}}}(w)\leq A_{2},\quad w\in\mathbb{D},$
    for some constants $A_{1}$ and $A_{2}$.
@@
$G(w)F(w)=I_{\mathbb{C}^{m}},\quad w\in\mathbb{D}.$
@@
Considering the composite of the unitary operator $V$ from Lemma 3.2 with $X$, we see that the invertible operator $X^{-1}V$ satisfies $(X^{-1}V)\mathcal{E}_{T}(w)=\mathcal{E}_{S^{*}_{n,\mathbb{C}^{m}}}(w)$ for all $w\in\mathbb{D}$. It follows that $T$ is similar to $S^{*}_{n,\mathbb{C}^{m}}$.